If a function f is not defined at x = a then the limit lim f(x) as x approaches a never exists. true or false


1
Expert's answer
2021-11-30T18:24:13-0500

The statement is False.

Counterexample

Let "f(x)=\dfrac{\sin x}{x}."

The function "f(x)" is not defined at "x=0."


"\lim\limits_{x\to 0^-}f(x)=\lim\limits_{x\to 0^-}\dfrac{\sin x}{x}=1"

"\lim\limits_{x\to 0^+}f(x)=\lim\limits_{x\to 0^+}\dfrac{\sin x}{x}=1"

We see that


"\lim\limits_{x\to 0^-}f(x)=1=\lim\limits_{x\to 0^+}f(x)"

Then "\lim\limits_{x\to 0}f(x)" exists, and


"\lim\limits_{x\to 0}f(x)=1"

The function "f(x)=\dfrac{\sin x}{x}" has a removable discontinuity at "x=0."


Therefore the statement "If a function "f" is not defined at "x = a" then the limit "\lim\limits_{x\to a}f(x)" never exists" is false.